IF tan(k+a)/p=tan(k+b)/q=tan(k+c)/r ,then prove that: ((p+q)/(p-q))*sin 2 (a-b) +((q-r)/(q+r))*sin 2 (b-c)+((r+p)(r-p))*sin 2 (c-a)=0

IF   tan(k+a)/p=tan(k+b)/q=tan(k+c)/r     ,then    prove that: ((p+q)/(p-q))*sin2(a-b) +((q-r)/(q+r))*sin2(b-c)+((r+p)(r-p))*sin2(c-a)=0

Grade:11

1 Answers

Rahul askIITiansExpert.IITR
13 years ago
the point here is to note that there is a certain rhythm in expressions

now

a/b=c/d

then, (a+b)/(a-b)=(c+d)/(c-d)

use this three times and write the expressiosn in the for form of tans and p,q etc

now put tan=sin/cos and transform

the procedure is length but will get you an expression which will be like sin2(a-b), since u know what the expanded form of sin(a-b) is.

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

DO APPROVE IF THE ANSWER HELPS YOU.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Regards,

Askiitians Experts

Rahul- IIT Roorkee